# IEEEexample:BSTcontrol

### v1

@@ 5 Discussion
There are a few key results to note from the outputs of the models, as shown in Tables 1-4. Firstly, despite their relative complexity and consistently high performance across the training regime, both the GraphSAGE and GUNet models achieve relatively high errors in the validation training set, regularly over double that of PointNet and MLP. One likely reason for such behaviour is that both GraphSAGE and GUNet are over-fitting to the data in training; this means the models are memorising the data they are being trained on, instead of learning the patterns or trends within the data.
Table 1 shows very little correlation across the results, with all models scoring the lowest RMSE for at least one metric. This clear discrepancy when compared to the rest of the Tables indicates that 5 Aerofoils is insufficient to reliably predict anything.
@@
Across the overall validation metrics, MLP predicts the values at the surface and the fluid the most accurately six out of eight times, in Tables 1, 2, and 4. Interestingly, in Table 2 despite being slightly less accurate at the surface and the fluid, PointNet outputs a more accurate $C_{L}$. Tables 3 and 4, however, show the expected results, with the model most accurately able to predict the vorticity at the surface, resulting in the most accurate $C_{L}$.
As evidenced in Figure 9, and Tables 1-4, it is only MLP that becomes consistently more accurate as the number of aerofoils increases. Interestingly, the point of lowest RMSE, across all experiments, is for PointNet using 20 aerofoils, despite the RMSE for $\omega$ being higher than that of MLP in Table 2. Notably, in Table 2, MLP has the lowest RMSE for every metric, aside for $C_{L}$, a pattern shared, with the overall RMSEs of Table 1. The fact that this pattern disappears as the number of aerofoils is increased could indicate that there is still a lack of sufficient training data required to provide consistent results. Further investigation is required to verify the cause of these inconsistencies.
@@
GraphSAGE and GUNet both experience large increases to the validation RMSE during the 55 aerofoil experiments, despite seeing a consistent decrease in test RMSE from experiment to experiment. These results combined with the longer inference required due to the more complicated nature of the models (as shown in Table 5) indicate a lack of suitability of GUNet and GraphSAGE for this problem statement. Whilst they train well, all evidence points to the models over-fitting to the training data.
@@
The foil-to-foil comparison, shown in Figure 4, provided further insight into the effect that the physical properties of each aerofoil had on the predictions of the models. The geometries of the most and least accurate four foils can be seen in Figures 5 and 6, respectively. When the validation aerofoils were compared to each other, four foils clearly stood out as having the highest errors across the models, whilst there were a large number of foils that resulted in relatively low RMSE.
@@
The differences between the two set of aerofoils are clear to see, the foils that performed better are thinner and have less of a camber. When comparing the RMSEs at each AoA for the less accurate foils, as shown in Figure 7, there is a visible correlation between the shape of the foils and whether the model was more accurate at low or high AoAs. The foils that were more rounded, with a larger lower camber were more accurate at lower angles of attack, whilst the foils with a larger mean camber were more accurate for higher angles of attack. Whereas the blades with a lower average RMSE for $C_{L}$ in Figure 8 are all less thick. Interestingly, the better performing blades all have points of lowest RMSE between 10 and 15 before having their highest RMSE at 20.
In Zhang et al. [30] found that a ‘drastically varying surface curvature in the leading edge’ [30] caused early flow detachments along the surface of the aerofoil. This separation causes the flow to reverse direction [30] which is more difficult for the models to predict than a simple and cohesive flow behaviour.
@@
A limitation of this paper was the selection of the aerofoils used in training; to allow the models to be compared, they were all trained on the same data, i.e., the first $nth$ foils in the dataset. It is usually good practice to randomise the dataset to reduce errors from training on too similar data. This could be further reason for the inconsistencies found in Tables 1 and 2.
@@
As stated earlier in Section 2, Bonnet et al. [2] provided a benchmark to data similar to that used for this paper, with the main differences being the variables calculated. Bonnet et al. also used the models to predict $C_{L}$ and $C_{D}$ directly, whereas this paper calculated them using the outputs. Whilst the results are comparable for the fluid predictions, the largest difference in accuracy lies in the predictions of the coefficients. One of the biggest drawbacks to using the panel method for calculating the coefficients is that it cannot be used to calculate the drag force acting on the aerofoil. Therefore, in order to use windAI_bench to find $C_{L}$ and $C_{D}$, the coefficients must be predicted using the models. However, if less data is incorporated into the models’ training, the predictions of $C_{L}$ and $C_{D}$ would be less accurate.
@@ 6 Conclusion
This paper found that whilst GraphSAGE and GUNet performed well during the testing phase, they performed poorly during the validation, meaning their ability to generalise to unseen data was inadequate relative to MLP and PointNet. Whilst these results for the latter models were relatively good, MLP showed far more consistency across the separate experiments, with the total RMSE found at surface, fluid, and $C_{L}$ decreasing with every increase of aerofoil number. To fully investigate the extent of this improving accuracy, research should be undertaken to verify the difference training on more than 150 aerofoils would make.
Whilst PointNet did show excellent results and would be a viable model to use with 55 foils (1375 simulations), other techniques for ensuring consistency would have to be investigated.

### v2

@@ 5 Discussion
There are a few key results to note from the outputs of the models, as shown in Tables 1-4. Firstly, despite their relative complexity and consistently high performance across the training regime, both the GraphSAGE and GUNet models achieve relatively high errors in the test set, regularly over double that of PointNet and MLP. The likely reason for such behaviour is that both GraphSAGE and GUNet are over-fitting to the data in training; this means the models are memorising the data they are being trained on, instead of learning the patterns or trends within the data.
Table 1 shows notably less correlation across the results than the following tables, with MLP and PointNet performing best over an equal number of metrics at the surface. This discrepancy when compared to the rest of the Tables, as well as the large RMSE across testing, indicates that 5 Aerofoils is insufficient to reliably predict the behaviour of the fluid within an acceptable margin of error.
@@
Across the overall test metrics, MLP predicts the values at the surface and the fluid the most accurately seven out of eight times, coming in second to PointNet only in Table 1. Interestingly, in Table 3 despite being slightly less accurate at predicting $\omega$ at the surface, PointNet predicts a more accurate $C_{L}$. Tables 3 and 4, however, show the expected results, with the model most accurately able to predict the vorticity at the surface, resulting in the most accurate $C_{L}$.
As evidenced in Fig. 5, and Tables 1-4, none of the aerofoil numbers show a consistent decrease in error for the prediction of $C_{L}$. Interestingly, the point of lowest RMSE for $C_{L}$, across all experiments, is for PointNet using 55 aerofoils, despite the RMSE for $\omega$ being higher than that of MLP in Table 2. Notably, in Table 3, MLP has the lowest total RMSE at both the surface and in the fluid, aside for $C_{L}$. Furthermore, despite the test RMSEs at the surface of the foils decreasing between Tables 3 and 4, the error for the calculation of $C_{L}$ actually increases. Whilst the precise reasoning for this requires further research, the most likely reason for this is the simplified nature of the panel method as it relies solely on $\omega$ at the surface and the flow conditions at infinity. Furthermore, the error for $\omega$ is the average over the whole of the surface of the foil, and larger errors at specific positions could have a larger effect of the overall calculation of $C_{L}$.
@@
GraphSAGE and GUNet both show notably higher errors in both the prediction of variables at the foil surface as well as in the calculations of $C_{L}$. As can be seen in Fig. 6, in comparison with MLP and PointNet which show a decrease in surface RMSE with every increase in foil number, the graph methods show increases in error at 150 foils. These results combined with the longer inference required due to the more complicated nature of the models (as shown in Table 5) indicate a lack of suitability of GUNet and GraphSAGE for this problem statement. Whilst they train well, all evidence points to the models over-fitting to the training data.
@@
The foil-to-foil comparison, shown in Fig. 4, provided further insight into the effect that the physical properties of each aerofoil had on the predictions of the models. The geometries of the most and least accurate four foils can be seen in Figs. 7 and 8, respectively. When the test aerofoils were compared to each other, four foils clearly stood out as having the highest errors across the models, whilst there were a large number of foils that resulted in relatively low RMSE.
@@
The differences between the two set of aerofoils are clear to see, the foils that performed better are thinner and have less of a camber. When comparing the RMSEs at each AoA for the less accurate foils, as shown in Fig. 9, there is a visible correlation between the shape of the foils and whether the model was more accurate at low or high AoAs. The foils that were more rounded, with a larger lower camber were more accurate at lower angles of attack, whilst the foils with a larger mean camber were more accurate for higher angles of attack. Whereas the blades with a lower average RMSE for $C_{L}$ in Fig. 10 are all less thick. Interestingly, the better performing blades all have points of lowest RMSE between 10 and 15 before having their highest RMSE at 20.
Zhang et al. [30] found that a ‘drastically varying surface curvature in the leading edge’ caused early flow detachments along the surface of the aerofoil. This separation causes the flow to reverse direction [30] which is more difficult for the models to predict than a simple and cohesive flow behaviour.
@@
As stated in Section 2, Bonnet et al. [9] provided a benchmark to data similar to that used for this paper, with the main differences being the variables calculated. Bonnet et al. also used the models to predict $C_{L}$ and $C_{D}$ directly, whereas this paper calculated them using the predicted outputs. Whilst the results are comparable for the fluid predictions, the largest difference in accuracy lies in the predictions of the coefficients. One of the biggest drawbacks to using the panel method for calculating the coefficients is that it cannot be used to calculate the drag force acting on the aerofoil. Therefore, in order to use windAI_bench to find $C_{L}$ and $C_{D}$, the coefficients must be predicted using the models.
The primary issue with the comparison of the two benchmarks is the lack of testing regime present in the AirfRANS paper. As this paper is more interested in the ability of these models to train on the data and then use that learnt knowledge to predict over unseen blades. To compare the training results of both papers would not provide any useful insights into the performance of the models and, as such, said comparison has not been performed.
@@ 6 Conclusion
This paper found that whilst GraphSAGE and GUNet performed well during the training phase, they performed poorly during testing, meaning their ability to generalise to unseen data was inadequate relative to MLP and PointNet. Whilst these results for the latter models were relatively good, MLP showed far more accuracy and consistency across the separate experiments, with the total RMSE found at surface, fluid, and $C_{L}$ decreasing with every increase of aerofoil number. To fully investigate the extent of this improving accuracy, research should be undertaken to verify the difference training on more than 150 aerofoils would make.